OpenStudy (anonymous):

How do you solve the problem -8+2x=-3xminus63 any help

OpenStudy (wasiqss):

take x at one side, like 3x-2x=63-8, then solve ahead

OpenStudy (anonymous):

is it -3x-63?

OpenStudy (wasiqss):

ohh sorry dint c that

OpenStudy (anonymous):

\[8+2x=-3x-63\]??

OpenStudy (anonymous):

add 3x to both sides, subtract 8 from both sides divide both sides by 5

OpenStudy (anonymous):

Always try to get rid of negative numbers. Move -8 to the right side and change the sign (+8). So on the right is -63 + 8. Move the -3x to the left and change the sign, yielding +3x + 2x. Now divide for the value of x (hint: it will be a negative number)

OpenStudy (anonymous):

x=-11

OpenStudy (anonymous):

mistinar is right and i was wrong. it says \[ -8+2x=-3x-63\] therefore 1) add 3x 2) add 8 3) divide by 5

OpenStudy (anonymous):

my ans is rite